Solve by using quadratic formula:
x^2 +4x+3 = 0
Thanks !!!

Answer by RAY100(1637)

x^2 +4x +3 =0
.
a=1,,,b=4,,,,c=3,,,,,from ax^2 +bx+c=0
.
x=[ -(b) +/-sqrt{(b)^2 - 4 (a)(c) }]/2(a)
.
x=[ -(4) +/- sqrt{ (4)^2 -4(1)(3)}]/2(1)
.
x=[ -4 +/sqrt{16-12}]/2
.
x= [ -4 +/- 2]/2
.
x= (-4-2)/2 = -3,,,,,x=(-4+2)/2 = -1
.
checking,,, factor original quadratic, x^2 +4x +3 = (x+3)(x+1)
.
set equal to zero ,,,x+3=0,,,x=-3,,,,,,x+1=0,,,x=-1,,,,,,,ok
.